OpenStudy (anonymous):

Write the Riemann sum to find the area under the graph of the function f(x) = x4 from x = 5 to x = 7.

OpenStudy (amistre64):

and what have you considered ....

OpenStudy (anonymous):

The area of the graph of the the function \[f(x)=x^4\] from x=5 to x=7 is the same as \[\int\limits_{5}^{7}x^4dx\] and know that \[\Delta x = \frac{ b-a }{ n} \] and\[x_{i} = a+i \Delta x\] \[\int\limits\limits_{5}^{7}x^4dx = \lim_{n \rightarrow \infty} \sum_{i=1}^{n}(5+\frac{ 2i }{ n })(\frac{ 2 }{ n })\] This is my first time answering a question so anyone feel free to correct me if I'm wrong